# Towards Multi-modal Graph Large Language Model

### v1

@@ Abstract
Multi-modal graphs, which integrate diverse multi-modal features and relations, are ubiquitous in real-world applications. However, existing multi-modal graph learning methods are typically trained from scratch for specific graph data and tasks, failing to generalize across various multi-modal graph data and tasks. To bridge this gap, we explore the potential of Multi-modal Graph Large Language Models (MG-LLM) to unify and generalize across diverse multi-modal graph data and tasks. We propose a unified framework of multi-modal graph data, task, and model, discovering the inherent multi-granularity and multi-scale characteristics in multi-modal graphs. Specifically, we present five key desired characteristics for MG-LLM: 1) unified space for multi-modal structures and attributes, 2) capability of handling diverse multi-modal graph tasks, 3) multi-modal graph in-context learning, 4) multi-modal graph interaction with natural language, and 5) multi-modal graph reasoning. We then elaborate on the key challenges, review related works, and highlight promising future research directions towards realizing these ambitious characteristics. Finally, we summarize existing multi-modal graph datasets pertinent for model training. We believe this paper can contribute to the ongoing advancement of the research towards MG-LLM for generalization across multi-modal graph data and tasks.
Index Terms:
@@
I Introduction
Multi-modal graphs, which integrate features from diverse modalities such as text, image, audio and video, as well as capture the complex intra-model and inter-modal relations, are becoming increasingly ubiquitous in real-world applications. From social networks [bhattacharyya2024heterogeneous] and e-commerce [wang2023fashionklip] platforms to scientific discovery in biomedicine and materials science [marini2024multimodal, ye2024construction], these complex data structures offer a richer, more holistic representation of interconnected entities than traditional unimodal graphs, which unlock more opportunities for advanced analytics, reasoning, and generation capabilities over multimodal information.
@@ I Introduction
The domain of multi-modal learning on graphs is rapidly advancing. We summarize the differences of this work with two most related papers as follows: 1) Ektefaie et al. [ektefaie2023multimodal] focus on the challenges and methodologies of combining modalities within graph-based machine learning, while we take a distinct perspective by envisioning MG-LLM as a unifying paradigm, leveraging the power of large language models to address the limitations of task-specific MGNNs. 2) Peng et al. [peng2024learningmultimodalgraphssurvey] provide a comprehensive survey of existing techniques, data types, and applications within multi-modal graph learning, while we focus on introducing a novel conceptual framework and put forth the vision of MG-LLM. By identifying the inherent characteristics of multi-modal graphs and the desired characteristics for MG-LLM, we chart a new research roadmap towards developing multi-modal graph large language models, thereby complementing surveys of the existing state-of-the-art by looking towards a next-generation paradigm.
II Towards a Unified View of Multi-modal Graph Data, Task, and Model
In this section, we introduce a unified framework of multi-modal graph data, task, and model, and remarks on the inherent characteristics in multi-modal graphs, serving as a foundation to discuss the desired characteristics of MG-LLM. The overall framework is shown in Figure 2.
II-A Unified Formulation of Multi-modal Graph Data
@@
Figure 2: Unified view of multimodal graph data, tasks and model towards MG-LLM.
@@
A modality is a distinct type or source of information associated with nodes or edges. Let =$\mathcal{M}=\{1,2,\dots,M\}$ denote the set of all modalities. For each $m\in\mathcal{M}$, we define a mapping $F_{m}$ from the modality-specific node feature space $\mathcal{V}$ to a shared representation space $\mathcal{X}$, such that $F_{m}:\mathcal{V}\to\mathcal{X}$. We define a mapping set =$\mathcal{F}=\{F_{m}\}_{m=1}^{M}$ for all $M$ modalities. The space $\mathcal{X}$ serves as a unified embedding space for all modalities. Similarly, we define a mapping ${F}_{m}$ from the modality-specific edge feature space $\mathcal{E}$ to the shared representation space $\mathcal{X}$, such that $\mathcal{F}_{m}:\mathcal{E}\to\mathcal{X}$. For simplicity, we reuse ${F}_{m}$ as the modality-$m$ map for both nodes and edges. We likewise leave out multimodal edges in the formulation, even though extending them would be straightforward. The features in different modalities could be texts, images, audios, videos, etc. A multi-modal graph can be defined by the quadruple $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{F},\mathcal{M})$.
@@
By instantiation of the modality set and feature mapping, we could obtain several classic types of multi-modal graphs which are ubiquitous in real-world applications [peng2024learningmultimodalgraphssurvey]. These types of multi-modal graphs share the same assumption that they can be decomposed by modality from different perspectives, i.e., feature, node, and graph:
  • Feature-level Multi-modal Graph, where the features of nodes or edges come from different modalities, i.e., $\mathcal{G}=\bigcup_{m=1}^{M}\mathcal{G}_{m}=\bigcup_{m=1}^{M}(\mathcal{V},\mathcal{E},\mathcal{F}_{m},\{m\}),$ where $m$ representing one modality in $M$ modalities. The feature of node $v$ could be represented as $\mathbf{x}(v)=\bigoplus_{m=1}^{M}\mathcal{F}_{m}(v)$. For example, on an e-commerce product graph [zhu2025mosaicmodalitiescomprehensivebenchmark], each node has the feature of product title and image, i.e., $\mathbf{x}(v)=\bigoplus(\mathcal{F}_{{\text{text}}}(v),\mathcal{F}_{{\text{image}}}(v))$.
  • Node-level Multi-modal Graph, where the nodes or edges come from different modalities, while each node or edge has unimodal features, i.e., $\mathcal{G}=\bigcup_{m=1}^{M}\mathcal{G}_{m}=\bigcup_{m=1}^{M}(\mathcal{V}_{m},\mathcal{E},\mathcal{F}_{m},\{m\}),$ where $m$ representing one modality in $M$ modalities, and $\mathcal{V}_{i}\bigcap\mathcal{V}_{j}=\emptyset$. For example, on a multimodal knowledge base [10.1007/978-3-030-41407-8_9], each node might be either an image or a textual description.
@@
Remark 1.
(Indecomposable Characteristics) Although practitioners may model their data with the aforementioned decomposable multi-modal graphs for convenience, most multi-modal graphs in real-world scenarios, with nodes and edges having features from various modalities, may not be easily decomposable to several uni-modal subgraphs. For instance, in a multi-modal graph where the text node says ‘The Transformer was incredible!’, the image node shows Optimus Prime (a central robot character from the Transformers movie series), and the knowledge node links to the movie Transformers, only joint reasoning over all three nodes can resolve the ambiguity and correctly interpret ‘Transformer’ as a film character rather than a neural network architecture. Due to the indecomposable characteristics, established multi-modal fusion techniques in other multi-modal fields [zhao2024deep] may fail to flexibly solve multi-modal graph problems, calling for the need of native modeling of multi-modal graph data in multi-modal graph large language models.
@@
  • a text sequence can be represented by a multi-modal graph with a single text-attributed node, i.e., $\mathcal{G}=\left(\mathcal{V}=\{v\},\mathcal{E}=\emptyset,\{F_{\text{text}}\},\{\text{text}\}\right)$, where $F_{\text{text}}$ is the text feature mapping function.
  • a text sequence, more granularly, can be represented by a multi-modal graph where each word is a node and sequential or semantic connections form edges, i.e., $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\{F_{\text{word}}\},\{\text{word}\}\right)$, where $\mathcal{V}=\{v_{1},\dots,v_{L}\},\mathcal{E}=\{(v_{i},v_{i+1})\}_{i=1}^{L-1}$, $L$ is the length of the text sequence, and $F_{\text{word}}$ is the word feature mapping function.
  • an image can be represented as a multi-modal graph where pixels are nodes and their grid-like inter-connections form edges, i.e., an $H\times W$ image can be $\mathcal{G}=\left(\mathcal{V}=\{v_{ij}\}_{i=1,j=1}^{H,W},\mathcal{E}_{\text{grid}},\{F_{\text{pixel}}\},\{\text{pixel}\}\right)$, where $\mathcal{E}_{\text{grid}}$ represents grid-like inter-connections, and $F_{\text{pixel}}$ is the pixel feature mapping function.
@@
  • A batch of images can be represented as a multi-modal graph with several image-attributed nodes without any edges, i.e., $\mathcal{G}=\left(\mathcal{V}=\{v_{1},\dots,v_{K}\},\mathcal{E}=\emptyset,\{F_{\text{image}}\},\{\text{image}\}\right)$, where $F_{\text{image}}$ is the image feature mapping function, and $K$ is the number of images.

### v2

# Towards multimodal graph large language model

@@ Abstract
Multi-modal graphs, which integrate diverse multi-modal features and relations, are ubiquitous in real-world applications. However, existing multi-modal graph learning methods are typically trained from scratch for specific graph data and tasks, failing to generalize across various multi-modal graph data and tasks. To bridge this gap, we explore the potential of Multi-modal Graph Large Language Models (MG-LLM) to unify and generalize across diverse multi-modal graph data and tasks. We propose a unified framework of multi-modal graph data, task, and model, discovering the inherent multi-granularity and multi-scale characteristics in multi-modal graphs. Specifically, we present five key desired characteristics for MG-LLM: 1) unified space for multi-modal structures and attributes, 2) capability of handling diverse multi-modal graph tasks, 3) multi-modal graph in-context learning, 4) multi-modal graph interaction with natural language, and 5) multi-modal graph reasoning. We then elaborate on the key challenges, review existing literature, and highlight promising future research directions towards realizing these ambitious characteristics. Finally, we summarize existing multi-modal graph datasets pertinent for model training. We believe this paper can contribute to the ongoing advancement of the research towards MG-LLM for generalization across multi-modal graph data and tasks.
keywords:
@@
1 Introduction
Multi-modal graphs, which integrate features from diverse modalities such as text, image, audio, and video, as well as capture the complex intra-model and inter-modal relations, are becoming increasingly ubiquitous in real-world applications. From social networks [1] and e-commerce [2] platforms to scientific discovery in biomedicine and materials science [3, 4], these complex data structures offer a richer, more holistic representation of interconnected entities than traditional unimodal graphs, which unlock more opportunities for advanced analytics, reasoning, and generation capabilities over multimodal information.
@@ 1 Introduction
The domain of multi-modal learning on graphs is rapidly advancing. We distinguish our work from several related lines of research. 1) Compared to surveys on multi-modal graph learning [5, 6], which primarily catalog existing techniques, our work introduces a novel conceptual framework and a forward-looking vision for a unified MG-LLM. 2) Unlike existing GraphLLMs [7, 8], which mainly focus on adapting unimodal graph data for LLMs, we address the more complex challenge of natively handling graphs with rich multi-modal attributes. 3) Distinct from general-purpose Omni-MLLMs [9], we argue for a specialized paradigm that deeply integrates graph-structured reasoning rather than treating graphs as just another input modality. By identifying the inherent characteristics of multi-modal graphs and the desired characteristics for MG-LLM, we chart a new research roadmap towards developing multi-modal graph large language models, thereby complementing surveys of the existing state-of-the-art by looking towards a next-generation paradigm.
Figure 2: Unified view of multimodal graph data, tasks, and model towards MG-LLM.
2 Towards a unified view of multi-modal graph data, task, and model
In this section, we introduce a unified framework of multi-modal graph data, task, and model, and remark on the inherent characteristics in multi-modal graphs, serving as a foundation to discuss the desired characteristics of MG-LLM. The overall framework is shown in Figure 2.
2.1 Unified formulation of multi-modal graph data
@@
A modality is a distinct type or source of information associated with nodes or edges. Let $\mathcal{M}=\{1,2,\dots,M\}$ denote the set of all modalities. We define a mapping set $\mathcal{F}=\{\mathcal{F}_{m}\}_{m=1}^{M}$ for all $M$ modalities. For each $m\in\mathcal{M}$, the mapping $\mathcal{F}_{m}$ maps from the modality-specific node feature space $\mathcal{V}$ to a shared representation space $\mathcal{X}$, such that $\mathcal{F}_{m}:\mathcal{V}\to\mathcal{X}$. The space $\mathcal{X}$ serves as a unified embedding space for all modalities. Similarly, we define a mapping $\mathcal{F}_{m}$ from the modality-specific edge feature space $\mathcal{E}$ to the shared representation space $\mathcal{X}$, such that $\mathcal{F}_{m}:\mathcal{E}\to\mathcal{X}$. For simplicity, we reuse $\mathcal{F}_{m}$ as the modality-$m$ map for both nodes and edges. We likewise leave out multimodal edges in the formulation, even though extending them would be straightforward. The features in different modalities could be texts, images, audios, videos, etc. A multi-modal graph can be defined by the quadruple $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{F},\mathcal{M})$. To make the formulation more concrete, a running example is shown in Figure 3.
Figure 3: An illustrative diagram and a running example of a multi-modal graph, take a small product graph with text and image nodes as an example.
@@
By instantiation of the modality set and feature mapping, we could obtain several classic types of multi-modal graphs which are ubiquitous in real-world applications [6]. These types of multi-modal graphs share the same assumption that they can be decomposed by modality from different perspectives, i.e., feature, node, and graph:
  • Feature-level Multi-modal Graph, where the features of nodes or edges come from different modalities, i.e., $\mathcal{G}=\bigcup_{m=1}^{M}\mathcal{G}_{m}=\bigcup_{m=1}^{M}(\mathcal{V},\mathcal{E},\mathcal{F}_{m},\{m\}),$ where $m$ representing one modality in $M$ modalities. The feature of node $v$ could be represented as $\mathbf{x}(v)=\bigoplus_{m=1}^{M}\mathcal{F}_{m}(v)$. For example, on an e-commerce product graph [10], each node has the feature of product title and image, i.e., $\mathbf{x}(v)=\bigoplus(\mathcal{F}_{{\text{text}}}(v),\mathcal{F}_{{\text{image}}}(v))$.
  • Node-level Multi-modal Graph, where the nodes or edges come from different modalities, while each node or edge has unimodal features, i.e., $\mathcal{G}=\bigcup_{m=1}^{M}\mathcal{G}_{m}=\bigcup_{m=1}^{M}(\mathcal{V}_{m},\mathcal{E},\mathcal{F}_{m},\{m\}),$ where $m$ representing one modality in $M$ modalities, and $\mathcal{V}_{i}\bigcap\mathcal{V}_{j}=\emptyset$. For example, on a multimodal knowledge base [11], each node might be either an image or a textual description.
@@
[…]
(Indecomposable characteristics) Although practitioners may model their data with the aforementioned decomposable multi-modal graphs for convenience, most multi-modal graphs in real-world scenarios, with nodes and edges having features from various modalities, may not be easily decomposable to several uni-modal subgraphs. For instance, in a multi-modal graph where the text node says ‘The Transformer was incredible!’, the image node shows Optimus Prime (a central robot character from the Transformers movie series), and the knowledge node links to the movie Transformers, only joint reasoning over all three nodes can resolve the ambiguity and correctly interpret ‘Transformer’ as a film character rather than a neural network architecture. Due to the indecomposable characteristics, established multi-modal fusion techniques in other multi-modal fields [13] may fail to flexibly solve multi-modal graph problems, calling for the need of native modeling of multi-modal graph data in multi-modal graph large language models.
@@
  • a text sequence can be represented by a multi-modal graph with a single text-attributed node, i.e., $\mathcal{G}=\left(\mathcal{V}=\{v\},\mathcal{E}=\emptyset,\{\mathcal{F}_{\text{text}}\},\{\text{text}\}\right)$, where $\mathcal{F}_{\text{text}}$ is the text feature mapping function.
  • a text sequence, more granularly, can be represented by a multi-modal graph where each word is a node and sequential or semantic connections form edges, i.e., $\mathcal{G}=\left(\mathcal{V},\mathcal{E},\{\mathcal{F}_{\text{word}}\},\{\text{word}\}\right)$, where $\mathcal{V}=\{v_{1},\dots,v_{L}\},\mathcal{E}=\{(v_{i},v_{i+1})\}_{i=1}^{L-1}$, $L$ is the length of the text sequence, and $\mathcal{F}_{\text{word}}$ is the word feature mapping function.
  • an image can be represented as a multi-modal graph where pixels are nodes and their grid-like inter-connections form edges, i.e., an $H\times W$ image can be $\mathcal{G}=\left(\mathcal{V}=\{v_{ij}\}_{i=1,j=1}^{H,W},\mathcal{E}_{\text{grid}},\{\mathcal{F}_{\text{pixel}}\},\{\text{pixel}\}\right)$, where $\mathcal{E}_{\text{grid}}$ represents grid-like inter-connections, and $\mathcal{F}_{\text{pixel}}$ is the pixel feature mapping function.
@@
  • A batch of images can be represented as a multi-modal graph with several image-attributed nodes without any edges, i.e., $\mathcal{G}=\left(\mathcal{V}=\{v_{1},\dots,v_{K}\},\mathcal{E}=\emptyset,\{\mathcal{F}_{\text{image}}\},\{\text{image}\}\right)$, where $\mathcal{F}_{\text{image}}$ is the image feature mapping function, and $K$ is the number of images.
  • An image-captioning dataset can be represented as a multi-modal graph where edges connect image-attributed nodes to their corresponding text-attributed caption nodes, i.e., for $K$ image-caption pairs, $\mathcal{G}=\left(\mathcal{V}_{\text{image}}\cup\mathcal{V}_{\text{text}},\mathcal{E}_{\text{image-text}},\{\mathcal{F}_{\text{image}},\mathcal{F}_{\text{text}}\},\{\text{image, text}\}\right)$, where $\mathcal{V}_{\text{image}}=\{v_{1},\dots,v_{K}\}$ are image nodes, $\mathcal{V}_{\text{text}}=\{u_{1},\dots,u_{K}\}$ are caption nodes, $\mathcal{E}_{\text{image-text}}=\{(v_{k},u_{k})\mid 1\leq k\leq K\}$ are edges connecting images to their captions, and $\mathcal{F}_{\text{image}}$ and $\mathcal{F}_{\text{text}}$ are the image and text feature mapping functions, respectively.